OpenStudy (anonymous):

FAN AND MEDAL Choose the correct simplification of the expression -3x2(5x - 4x2 - 6).

OpenStudy (anonymous):

-7x4 + 2x3 - 9x2 12x4 - 15x3 + 18x2 12x4 + 15x3 + 18x2 -12x4 + 15x3 - 18x2

OpenStudy (anonymous):

b or c i think

OpenStudy (solomonzelman):

When you multiply exponents, you add them. \(\large\color{blue}{ \bf m^3 \times m^{4}=m^{3+4}=m^{7} }\) you are getting the numbers in terms of absolute values right, but you are getting wrong the negatives and positives.

OpenStudy (anonymous):

ahh so its b

OpenStudy (anonymous):

?

OpenStudy (solomonzelman):

yes... that's right !

OpenStudy (anonymous):

thankyou

OpenStudy (solomonzelman):

Anytime !